# Galois Correspondence for Partial Groupoid Actions

### v1

@@ Abstract.
We prove a Galois correspondence theorem for groupoids acting orthogonally and partially on commutative rings. We also consider partial actions that are not orthogonal, presenting two correspondences in this case: one for strongly Galois partial groupoid actions and one for global groupoid actions (without restriction). Some examples are illustrated.
@@ 1. Introduction
So, the first objective of this work is to extend the Galois correspondence presented in [4] to a partial groupoid action without asking the group-type property. For this, we will relate a partial Galois extension to the Galois extension regarding its globalization. The partial-to-global relation will give us a more general partial Galois correspondence, as we will present in Section 3.
The second objective is to improve the assumption of Galois correspondence theorems for groupoid actions in general. In fact, all the Galois correspondences that are found in the literature asks that the action must be orthogonal, that is, $A=\bigoplus_{e\in\mathcal{G}_{0}}A_{e}$, where $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a partial action of a groupoid $\mathcal{G}$ on a ring $A$ and $\mathcal{G}_{0}$ is the set of identities of $\mathcal{G}$. This hypothesis covers a broad class of actions, but is still restrictive because it does not cover them all. Every action can be orthogonalized, but the problem is that if we have a Galois extension $A|_{A^{\varepsilon}}$ regarding an orthogonalization $\varepsilon$, it does not implies that the extension $A|_{A^{\alpha}}$ is Galois regarding the original action $\alpha$. In Section 4 we present some examples, as well as we prove two Galois correspondences without the orthogonalization hypothesis.
We give now a brief outline of the paper. In the second section we present basic definitions and results about partial groupoid actions and Galois theory. In the third section we prove our first main result, which states that there is a one-to-one correspondence between the wide subgroupoids $\mathcal{H}$ of the groupoid $\mathcal{G}$ and determined separable, $\alpha$-strong $A^{\alpha}$-subalgebras $C$ of the algebra $A$, where $\alpha$ is a partial action of $\mathcal{G}$ on $A$ (Theorem 3.6). In Section 4 we define orthogonalization of a given partial action and we study which properties of Galois theory remain valid under orthogonalization. We also define strongly $\alpha$-Galois extensions and we present in Theorem 4.11 a Galois correspondence in this case, which is our second main result. Some examples are illustrated. We end the paper with our third main result, which is a Galois correspondence for global actions without restrictions on the action (Theorem 4.16).
@@ 2. Preliminaries
We recall that a groupoid $\mathcal{G}$ is a small category in which every morphism is an isomorphism. We denote by $\mathcal{G}_{0}$ the set of objects of $\mathcal{G}$. Observe that $\text{id}:\mathcal{G}_{0}\rightarrow\mathcal{G}$, given by $\text{id}(x)=\text{id}_{x}$, is an injective map and whence we identify $\mathcal{G}_{0}\subseteq\mathcal{G}$. Given $g\in\mathcal{G}$, the domain and the range of $g$ will be denoted by $d(g)$ and $r(g)$, respectively. Hence, $d(g)=g^{-1}g$ and $r(g)=gg^{-1}$. For all $g,h\in\mathcal{G}$, we write $\exists gh$ whenever the product $gh$ is defined. We denote $\mathcal{G}_{2}=\{(g,h)\in\mathcal{G}\times\mathcal{G}:\exists gh\}$.
@@
Given a nonempty set $S$, we recall that the coarse groupoid associated to $S$ is the groupoid $S^{2}=S\times S$, where $\exists(s,t)(u,v)$ if and only if $s=v$ and in this case $(s,t)(u,v)=(u,t)$. If $\mathcal{G}$ is a connected groupoid, we have $\mathcal{G}\simeq\mathcal{G}_{0}^{2}\times\mathcal{G}(e)$ (c.f. [13, Proposition 3.3.6]), and this isomorphism does not depend on the choice of $e\in\mathcal{G}_{0}$. Notice that every two finite coarse groupoids with the same amount of identities are isomorphic. Hence we will just denote by $\mathcal{A}_{n}$ the coarse groupoid with $n$ identities. Thus, if $\mathcal{G}$ is a finite connected groupoid with cardinality $|\mathcal{G}_{0}|=n$ and $e\in\mathcal{G}_{0}$, we have $\mathcal{G}\simeq\mathcal{A}_{n}\times\mathcal{G}(e)$.
@@ 2.1. Partial Groupoid Actions
The first thing to note is that condition (P1) is not the same as reference [2], because we ask that $A=\sum_{e\in\mathcal{G}_{0}}A_{e}$, and in [2] this condition was not requested. When $\mathcal{G}$ is a group with identity element $1_{\mathcal{G}}$, this condition translates into $A=A_{1_{\mathcal{G}}}$. If we don’t ask for this, it may happen that a partial groupoid action is not a partial group action, even when the groupoid is a group.
@@
We say that a partial action $\alpha$ is preunital if $A_{e}$ is a unital ring with identity $1_{e}$ central and idempotent, so that $A_{e}=A1_{e}=1_{e}A$, for all $e\in\mathcal{G}_{0}$. We say that a partial action $\alpha$ is unital if $A_{g}$ is a unital ring with identity $1_{g}$ central and idempotent, so that $A_{g}=A1_{g}=1_{g}A$, for all $g\in\mathcal{G}$.
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a partial action of a groupoid $\mathcal{G}$ on a ring $A$. A global action $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $B$ is said to be a globalization of $\alpha$ if, for all $e\in\mathcal{G}_{0}$, there is $\varphi_{e}:A_{e}\to B_{e}$ monomorphism of rings such that
@@ 2.2. Galois Theory
In this subsection we present some basic concepts regarding Galois theory.
Definition 2.5.
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a partial action of a groupoid $\mathcal{G}$ on a ring $A$. We define
@@
Let $\mathcal{G}$ be a groupoid, $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a partial action of $\mathcal{G}$ on a ring $A$ and $\mathcal{H}$ be a subgroupoid of $\mathcal{G}$. We define the restriction of $\alpha$ to $\mathcal{H}$ as
@@
Let $\mathcal{G}$ be a groupoid, $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ a partial action of $\mathcal{G}$ on a ring $A$ and $C$ a $A^{\alpha}$-subalgebra of $A$. We define
@@
Let $C$ be an $A^{\alpha}$-subalgebra of a ring $A$, where $\alpha$ is a partial action of a groupoid $\mathcal{G}$ on $A$. We say that $C$ is:
@@ 3. Correspondence Theorem
This section is intended to prove a Galois correspondence for partial groupoid actions.
Let $e\in\mathcal{G}_{0}$. Consider $\mathcal{G}(-,e)=\{g\in\mathcal{G}:r(g)=e\}$. Suppose that $\mathcal{G}(-,e)=\{e=g_{1,e},g_{2,e}\ldots,g_{n_{e},e}\}$, where $n_{e}$ is the cardinality $|\mathcal{G}(-,e)|$ of $\mathcal{G}(-,e)$. Consider the following assertion:
@@
In [2, Theorem 5.1], it was proved that if (1) holds, then $A$ being a partial $\alpha$-Galois extension of $A^{\alpha}$ implies that $B$ is a $\beta$-Galois extension of $B^{\beta}$. We shall prove that this hypothesis is superfluous. So we start this section by improving [2, Theorem 5.1].
@@
Let $\alpha$ be a unital, orthogonal partial action of a finite groupoid $\mathcal{G}$ on a ring. Consider $\beta$ globalization of $\alpha$, global action of $\mathcal{G}$ on a ring $B$. The following statements are equivalent:
@@
From now on, consider $\alpha$ an orthogonal partial action of a finite groupoid $\mathcal{G}$ on a commutative ring $A$. We fix the notation $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ to denote the globalization of $\alpha$, which is a global action on a ring $B$. According to [2, Theorem 2.1], it follows that $B=\prod_{e\in\mathcal{G}_{0}}B_{e}$, so that $\beta$ is an orthogonal action. To shorten the notation, we will assume $A_{e}\subseteq B_{e}$, for all $e\in\mathcal{G}_{0}$, instead of $\varphi_{e}(A_{e})\subseteq B_{e}$. We will also denote by $1_{g}$ the identity of $A_{g}$ and by $1_{g}^{\prime}$ the identity of $B_{g}$, for all $g\in\mathcal{G}$.
@@
Now we prove that $C$ is $\alpha$-strong. Let $g,h\in\mathcal{G}$ be such that $\exists g^{-1}h$, $g^{-1}h\notin\mathcal{G}_{C}$ and $e\in A_{g}\cup A_{h}$ a nonzero idempotent. Since $B^{\beta|_{\mathcal{H}}}$ is $\beta$-strong and $g^{-1}h\notin\mathcal{G}_{C}\supseteq\mathcal{H}=\mathcal{G}_{B^{\beta|_{\mathcal{H}}}}^{\prime}$, if $e1_{g}1_{h}\neq 0$, there is $b\in B^{\beta|_{\mathcal{H}}}$ such that $\beta_{g}(b1_{g^{-1}}^{\prime})e1_{g}1_{h}\neq\beta_{h}(b1_{h^{-1}}^{\prime})e1_{g}1_{h}$. Hence $b1_{A}\in B^{\beta|_{\mathcal{H}}}1_{A}=C$ and $\alpha_{g}(b1_{A}1_{g^{-1}})e1_{g}1_{h}=\alpha_{g}(b1_{g^{-1}})e1_{g}1_{h}=\beta_{g}(b1_{g^{-1}}^{\prime})e1_{g}1_{h}\neq\beta_{h}(b1_{h^{-1}}^{\prime})e1_{g}1_{h}=\alpha_{h}(b1_{h^{-1}})e1_{g}1_{h}=\alpha_{h}(b1_{A}1_{h}^{-1})e1_{g}1_{h}$, from where we obtain $\alpha_{g}(b1_{A}1_{g^{-1}})e\neq\alpha_{h}(b1_{A}1_{h^{-1}})e$. If $e1_{g}1_{h}=0$ and $e\in A_{g}$, then $\alpha_{g}(1_{A}1_{g^{-1}})e=1_{g}e=e\neq 0=e1_{g}1_{h}=e1_{h}=\alpha_{h}(1_{A}1_{h^{-1}})e$. The case such that $e\in A_{h}$ is similar.
Finally, it only remains to us to show that $\mathcal{G}_{C}\subseteq\mathcal{H}$, since we already have that $\mathcal{H}\subseteq\mathcal{G}_{C}$. Assume, by contradiction, that there is $g\in\mathcal{G}_{C}\setminus\mathcal{H}$. Since $B^{\beta|_{\mathcal{H}}}$ is $\beta$-strong, there is $b\in B^{\beta|_{\mathcal{H}}}$ such that $\beta_{g}(b1_{g^{-1}}^{\prime})1_{g}\neq\beta_{r(g)}(b1_{r(g)}^{\prime})1_{g}=b1_{g}$. Hence, $\alpha_{g}(b1_{A}1_{g^{-1}}^{\prime})=\beta_{g}(b1_{g^{-1}}^{\prime})1_{g}\neq b1_{g}=b1_{A}1_{g}$, and it is a contradiction because $g\in\mathcal{G}_{C}$ and $b1_{A}\in B^{\beta|_{\mathcal{H}}}1_{A}=C$. Thus $\mathcal{H}$ must coincide with $\mathcal{G}_{C}$, ending the proof. ∎
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a unital, orthogonal partial action of a groupoid $\mathcal{G}$ on a ring $A$. Then $A|_{A^{\alpha}}$ is a partial $\alpha$-Galois extension if and only if $A$ is $A^{\alpha}$-separable and $\alpha$-strong.
@@
By the discussion above, it only remains to us to prove that if $A$ is $A^{\alpha}$-separable and $\alpha$-strong, then $A$ is partial $\alpha$-Galois over $A^{\alpha}$.
For that, let $e=\sum_{i=1}^{n}x_{i}\otimes y_{i}\in A\otimes_{A^{\alpha}}A$ the idempotent of separability of $A$. That is, $\mu(e)=\sum_{i=1}^{n}x_{i}y_{i}=1_{A}$ and $(a\otimes 1_{A}-1_{A}\otimes a)e=0$, for all $a\in A$, where $\mu:A\otimes_{A^{\alpha}}A$ is the multiplication map. For all $g\in\mathcal{G}$, define $e_{g}=m((\text{Id}_{A}\otimes\alpha_{g}(-1_{g^{-1}}))(e))\in A_{g}$. We have that $\text{Id}_{A}\otimes\alpha_{g}(-1_{g^{-1}})$ is an endomorphism of $A\otimes_{A^{\alpha}}A$. Besides that, $\mu$ is a ring homomorphism, since $A$ is commutative. Hence, for all $g\in\mathcal{G}$ we have
@@
By [2, Theorem 4.6], there is a ring homomorphism $\psi_{\mathcal{G}_{C}}:A\to B^{\prime}$ such that $\psi_{\mathcal{G}_{C}}|_{A^{\alpha|_{\mathcal{G}_{C}}}}:A^{\alpha|_{\mathcal{G}_{C}}}\to(B^{\prime})^{\beta^{\prime}}$ is a ring homomorphism and $\psi_{\mathcal{G}_{C}}(a)1_{A}=a$, for all $a\in A^{\alpha|_{\mathcal{G}_{C}}}$. Let $C^{\prime}=\psi_{\mathcal{G}_{C}}(C)$.
@@
Let $x_{i}^{\prime}=\psi_{\mathcal{G}_{C}}(x_{i})$ and $y_{i}^{\prime}=\psi_{\mathcal{G}_{C}}(y_{i})$. Then $\sum_{i=1}^{n}x_{i}^{\prime}y_{i}^{\prime}=\psi_{\mathcal{G}_{C}}\left(\sum_{i=1}^{n}x_{i}y_{i}\right)=\psi_{\mathcal{G}_{C}}(1_{A})=1_{B^{\prime}}$. Consider now $e\in\mathcal{G}_{0}$ and $\{h_{1,e}=e,h_{2,e},\ldots,h_{m_{e},e}\}=\mathcal{G}_{C}(-,e)=\{g\in\mathcal{G}_{C}:r(g)=e\}$, where $m_{e}>0$. By [2, pg. 15], $1_{e}^{\prime}=v_{1,e}+\cdots+v_{m_{e},e}$, where $v_{1,e}=1_{e}$ and $v_{j,e}=(1_{e}^{\prime}-1_{e})(1_{e}^{\prime}-\beta_{h_{2,e}}(1_{d(h_{2,e})}))\cdots(1_{e}^{\prime}-\beta_{h_{j-1,e}}(1_{d(h_{j-1,e})}))\beta_{h_{j,e}}(1_{d(h_{j,e})})$.

### v2

@@ Abstract.
We prove a Galois correspondence theorem for groupoids acting orthogonally and partially on commutative rings. We also consider partial actions that are not orthogonal, presenting two correspondences in this case: one for strongly Galois partial groupoid actions and one for global groupoid actions (without restriction). Some examples are presented.
@@ 1. Introduction
So, the first objective of this work is to extend the Galois correspondence presented in [4] to a partial groupoid action without requiring the group-type property. For this, we will relate a partial Galois extension to the Galois extension regarding its globalization. The partial-to-global relation will give a more general partial Galois correspondence, as we will present in Section 3.
The second objective is to improve the assumption of Galois correspondence theorems for groupoid actions in general. In fact, all the Galois correspondences found in the literature require the action to be orthogonal, that is, $A=\bigoplus_{e\in\mathcal{G}_{0}}A_{e}$, where $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a partial action of a groupoid $\mathcal{G}$ on a ring $A$ and $\mathcal{G}_{0}$ is the set of identities of $\mathcal{G}$. This hypothesis covers a broad class of actions, but is still restrictive because it does not cover them all. Every action can be orthogonalized, but the problem is that if we have a Galois extension $A|_{A^{\varepsilon}}$ regarding an orthogonalization $\varepsilon$, it does not imply that the extension $A|_{A^{\alpha}}$ is Galois regarding the original action $\alpha$. In Section 4 we present some examples, as well as we prove two Galois correspondences without the orthogonal hypothesis.
We give now a brief outline of the paper. In the second section we present basic definitions and results about partial groupoid actions and Galois theory. In the third section we prove our first main result, which states that there is a one-to-one correspondence between the wide subgroupoids $\mathcal{H}$ of the groupoid $\mathcal{G}$ and determined separable, $\alpha$-strong $A^{\alpha}$-subalgebras $C$ of the algebra $A$, where $\alpha$ is a partial action of $\mathcal{G}$ on $A$ (Theorem 3.6). In Section 4 we define orthogonalization of a given partial action and we study which properties of Galois theory remain valid under orthogonalization. We also define strongly $\alpha$-Galois extensions and we present in Theorem 4.11 a Galois correspondence in this case, which is our second main result. Some examples are presented. We end the paper with our third main result, which is a Galois correspondence for global actions without restrictions on the action (Theorem 4.16).
By rings, we mean unital rings. We also use the notation $A\triangleleft B$ to indicate that $A$ is a two-sided ideal of the ring $B$.
@@ 2. Preliminaries
We recall that a groupoid $\mathcal{G}$ is a small category in which every morphism is invertible. Given a morphism $g\in\mathcal{G}$, we denote its domain and range by $d(g)$ and $r(g)$, respectively. Hence, $d(g)=g^{-1}g$ and $r(g)=gg^{-1}$. Furthermore, $\mathcal{G}_{0}$ denotes the set of objects of $\mathcal{G}$. Instead of using $Mor(\mathcal{G})$ to denote the set of morphisms of $\mathcal{G}$, we identify $\mathcal{G}$ with $Mor(\mathcal{G})$. We denote
$\mathcal{G}_{2}=\{(g,h)\in\mathcal{G}\times\mathcal{G}:\text{the product }gh\text{ is defined}\}.$
For each $x\in\mathcal{G}_{0}$, we identify $x$ with the identity morphism $id_{x}$, thus considering $\mathcal{G}_{0}\subseteq\mathcal{G}$.
@@
Given a nonempty set $S$, we recall that the coarse groupoid associated to $S$ is the groupoid $S^{2}=S\times S$, where the product $(s,t)(u,v)$ is defined if and only if $s=v$ and in this case $(s,t)(u,v)=(u,t)$. If $\mathcal{G}$ is a connected groupoid, we have $\mathcal{G}\simeq\mathcal{G}_{0}^{2}\times\mathcal{G}(e)$ (c.f. [13, Proposition 3.3.6]), and this isomorphism does not depend on the choice of $e\in\mathcal{G}_{0}$. Notice that every two finite coarse groupoids with the same amount of identities are isomorphic. Hence we will just denote by $\mathcal{A}_{n}$ the coarse groupoid with $n$ identities. Thus, if $\mathcal{G}$ is a finite connected groupoid with cardinality $|\mathcal{G}_{0}|=n$ and $e\in\mathcal{G}_{0}$, we have $\mathcal{G}\simeq\mathcal{A}_{n}\times\mathcal{G}(e)$.
@@ 2.1. Partial Groupoid Actions
The first thing to note is that condition (P1) is not the same as reference [2], because we require that $A=\sum_{e\in\mathcal{G}_{0}}A_{e}$, and in [2] this condition was not requested. When $\mathcal{G}$ is a group with identity element $1_{\mathcal{G}}$, this condition translates into $A=A_{1_{\mathcal{G}}}$. If we do not require for this, it may happen that a partial groupoid action is not a partial group action, even when the groupoid is a group.
@@
We say that the partial action $\alpha$ is preunital if each $A_{e}$ is generated by a central idempotent of $A$, denoted by $1_{e}$, for $e\in\mathcal{G}_{0}$. We say that a partial action $\alpha$ is unital if each $A_{g}$ is generated by a central idempotent $1_{g}$, for $g\in\mathcal{G}$.
@@
Let $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ be a partial action of a groupoid $\mathcal{G}$ on a ring $A$. A global action $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ of $\mathcal{G}$ on a ring $B$ is said to be a globalization of $\alpha$ if, for all $e\in\mathcal{G}_{0}$, there is a monomorphism of rings $\varphi_{e}:A_{e}\to B_{e}$ such that
@@ 2.2. Galois Theory
In this subsection we present some basic concepts regarding Galois theory. From now on, assume that $\mathcal{G}$ is a groupoid, and $\alpha=(A_{g},\alpha_{g})_{g\in\mathcal{G}}$ is a unital partial action of $\mathcal{G}$ on a ring $A$. We define
@@
Let $\mathcal{H}$ be a subgroupoid of $\mathcal{G}$. We define the restriction of $\alpha$ to $\mathcal{H}$ as
@@
Let $C$ be a $A^{\alpha}$-subalgebra of $A$. We define
@@
Let $C$ be a $A^{\alpha}$-subalgebra of a ring $A$. We say that $C$ is:
@@ 3. Correspondence Theorem
This section is intended to prove a Galois correspondence for partial groupoid actions. From now on, assume that the groupoid $\mathcal{G}$ is finite.
Take $e\in\mathcal{G}_{0}$. Consider $\mathcal{G}(-,e)=\{g\in\mathcal{G}:r(g)=e\}$. Suppose that $\mathcal{G}(-,e)=\{e=g_{1,e},g_{2,e}\ldots,g_{n_{e},e}\}$, where $n_{e}$ is the cardinality $|\mathcal{G}(-,e)|$ of $\mathcal{G}(-,e)$. Consider the following assertion:
@@
In [2, Theorem 5.1], it was proved that if (1) holds, then $A$ being a partial $\alpha$-Galois extension of $A^{\alpha}$ implies that $B$ is a $\beta$-Galois extension of $B^{\beta}$, where $\beta$ is the globalization of $\alpha$, a global action of $\mathcal{G}$ on a ring $B$. We shall prove that this hypothesis is superfluous. So we start this section by improving [2, Theorem 5.1].
@@
Assume that $\alpha$ is orthogonal. Consider a global action $\beta$ of $\mathcal{G}$ on a ring $B$, and assume that $\beta$ is a globalization of $\alpha$. The following statements are equivalent:
@@
Throughout this section, assume that $\alpha$ is orthogonal and that the ring $A$ is commutative. We fix the notation $\beta=(B_{g},\beta_{g})_{g\in\mathcal{G}}$ to denote a global action of $\mathcal{G}$ on a ring $B$ that is a globalization of $\alpha$. According to [2, Theorem 2.1], it follows that $B=\prod_{e\in\mathcal{G}_{0}}B_{e}$, so that $\beta$ is an orthogonal action. To shorten the notation, we will assume $A_{e}\subseteq B_{e}$, for all $e\in\mathcal{G}_{0}$, instead of $\varphi_{e}(A_{e})\subseteq B_{e}$. We will also denote by $1_{g}$ the identity of $A_{g}$ and by $1_{g}^{\prime}$ the identity of $B_{g}$, for all $g\in\mathcal{G}$.
@@
Now we prove that $C$ is $\alpha$-strong. Let $g,h\in\mathcal{G}$ be such that $(g^{-1},h)\in\mathcal{G}_{2}$, $g^{-1}h\notin\mathcal{G}_{C}$ and $e\in A_{g}\cup A_{h}$ a nonzero idempotent. Since $B^{\beta|_{\mathcal{H}}}$ is $\beta$-strong and $g^{-1}h\notin\mathcal{G}_{C}\supseteq\mathcal{H}=\mathcal{G}_{B^{\beta|_{\mathcal{H}}}}^{\prime}$, if $e1_{g}1_{h}\neq 0$, there is $b\in B^{\beta|_{\mathcal{H}}}$ such that $\beta_{g}(b1_{g^{-1}}^{\prime})e1_{g}1_{h}\neq\beta_{h}(b1_{h^{-1}}^{\prime})e1_{g}1_{h}$. Hence $b1_{A}\in B^{\beta|_{\mathcal{H}}}1_{A}=C$ and $\alpha_{g}(b1_{A}1_{g^{-1}})e1_{g}1_{h}=\alpha_{g}(b1_{g^{-1}})e1_{g}1_{h}=\beta_{g}(b1_{g^{-1}}^{\prime})e1_{g}1_{h}\neq\beta_{h}(b1_{h^{-1}}^{\prime})e1_{g}1_{h}=\alpha_{h}(b1_{h^{-1}})e1_{g}1_{h}=\alpha_{h}(b1_{A}1_{h}^{-1})e1_{g}1_{h}$, from where we obtain $\alpha_{g}(b1_{A}1_{g^{-1}})e\neq\alpha_{h}(b1_{A}1_{h^{-1}})e$. If $e1_{g}1_{h}=0$ and $e\in A_{g}$, then $\alpha_{g}(1_{A}1_{g^{-1}})e=1_{g}e=e\neq 0=e1_{g}1_{h}=e1_{h}=\alpha_{h}(1_{A}1_{h^{-1}})e$. The case $e\in A_{h}$ is similar.
Finally, it only remains to show that $\mathcal{G}_{C}\subseteq\mathcal{H}$, since we already have that $\mathcal{H}\subseteq\mathcal{G}_{C}$. Assume, by contradiction, that there is $g\in\mathcal{G}_{C}\setminus\mathcal{H}$. Since $B^{\beta|_{\mathcal{H}}}$ is $\beta$-strong, there is $b\in B^{\beta|_{\mathcal{H}}}$ such that $\beta_{g}(b1_{g^{-1}}^{\prime})1_{g}\neq\beta_{r(g)}(b1_{r(g)}^{\prime})1_{g}=b1_{g}$. Hence, $\alpha_{g}(b1_{A}1_{g^{-1}}^{\prime})=\beta_{g}(b1_{g^{-1}}^{\prime})1_{g}\neq b1_{g}=b1_{A}1_{g}$, and it is a contradiction because $g\in\mathcal{G}_{C}$ and $b1_{A}\in B^{\beta|_{\mathcal{H}}}1_{A}=C$. Thus $\mathcal{H}$ must coincide with $\mathcal{G}_{C}$, ending the proof. ∎
@@
$A|_{A^{\alpha}}$ is a partial $\alpha$-Galois extension if and only if $A$ is $A^{\alpha}$-separable and $\alpha$-strong.
@@
By the discussion above, it only remains to prove that if $A$ is $A^{\alpha}$-separable and $\alpha$-strong, then $A$ is partial $\alpha$-Galois over $A^{\alpha}$.
For that, let $e=\sum_{i=1}^{n}x_{i}\otimes y_{i}\in A\otimes_{A^{\alpha}}A$ be the idempotent of separability of $A$. That is, $\mu(e)=\sum_{i=1}^{n}x_{i}y_{i}=1_{A}$ and $(a\otimes 1_{A}-1_{A}\otimes a)e=0$, for all $a\in A$, where $\mu:A\otimes_{A^{\alpha}}A$ is the multiplication map. For all $g\in\mathcal{G}$, define $e_{g}=m((\text{Id}_{A}\otimes\alpha_{g}(-1_{g^{-1}}))(e))\in A_{g}$. We have that $\text{Id}_{A}\otimes\alpha_{g}(-1_{g^{-1}})$ is an endomorphism of $A\otimes_{A^{\alpha}}A$. Besides that, $\mu$ is a ring homomorphism, since $A$ is commutative. Hence, for all $g\in\mathcal{G}$ we have
@@
By [2, Theorem 4.6], there is a ring homomorphism $\psi_{\mathcal{G}_{C}}:A\to B^{\prime}$ such that $\psi_{\mathcal{G}_{C}}|_{A^{\alpha|_{\mathcal{G}_{C}}}}:A^{\alpha|_{\mathcal{G}_{C}}}\to(B^{\prime})^{\beta^{\prime}}$ is a ring homomorphism and $\psi_{\mathcal{G}_{C}}(a)1_{A}=a$, for all $a\in A^{\alpha|_{\mathcal{G}_{C}}}$. Consider $C^{\prime}=\psi_{\mathcal{G}_{C}}(C)$.
@@
Take $x_{i}^{\prime}=\psi_{\mathcal{G}_{C}}(x_{i})$ and $y_{i}^{\prime}=\psi_{\mathcal{G}_{C}}(y_{i})$. Then $\sum_{i=1}^{n}x_{i}^{\prime}y_{i}^{\prime}=\psi_{\mathcal{G}_{C}}\left(\sum_{i=1}^{n}x_{i}y_{i}\right)=\psi_{\mathcal{G}_{C}}(1_{A})=1_{B^{\prime}}$. Consider now $e\in\mathcal{G}_{0}$ and $\{h_{1,e}=e,h_{2,e},\ldots,h_{m_{e},e}\}=\mathcal{G}_{C}(-,e)=\{g\in\mathcal{G}_{C}:r(g)=e\}$, where $m_{e}>0$. By [2, pg. 15], $1_{e}^{\prime}=v_{1,e}+\cdots+v_{m_{e},e}$, where $v_{1,e}=1_{e}$ and $v_{j,e}=(1_{e}^{\prime}-1_{e})(1_{e}^{\prime}-\beta_{h_{2,e}}(1_{d(h_{2,e})}))\cdots(1_{e}^{\prime}-\beta_{h_{j-1,e}}(1_{d(h_{j-1,e})}))\beta_{h_{j,e}}(1_{d(h_{j,e})})$.